OpenStudy (anonymous):

Which value of x makes the following equation true? 4(x - 2) = -2x + 40 12 8 9 11

OpenStudy (the_fizicx99):

Same thing :> 4x - 8 = -2x + 40 +8 +8 4x = -2x + 48 +2x +2x 6x = 48 /6 /6 x = 8

OpenStudy (anonymous):

THANKS!!!!!!!!!!!!!:D :):D

OpenStudy (the_fizicx99):

yw